OpenStudy (anonymous):

(-2,-2) an ordered pair of y=3x-8

OpenStudy (anonymous):

(-2 , -2) means that f(-2) = -2 now plug in -2 and you'll get the result y = 3(-2) + 8 y = -6 + 8 y = -2

OpenStudy (anonymous):

yes it is an ordered pair for y